Are these true or false
square root of [ln(A)]=ln(A^(1/2))

2. log(square root of A)=(1/2)log(A)

3. ln(A)ln(B)=ln(A)+ln(B)

4. p⋅ln(A)=ln(A^p)

5. log(AB)=log(A)+log(B)

6. log(A)/log(B)=log(A)−log(B)

Found 2 solutions by t0hierry, ikleyn:
Answer by t0hierry(194)

Are these true or false
square root of [ln(A)]=ln(A^(1/2)) FALSE
2. log(square root of A)=(1/2)log(A) TRUE
3. ln(A)ln(B)=ln(A)+ln(B) FALSE it's ln(AB) = ln(A) + ln(B)
4. p⋅ln(A)=ln(A^p) TRUE
5. log(AB)=log(A)+log(B) TRUE
6. log(A)/log(B)=log(A)−log(B) FALSE it's ln(A/B) = ln(A) - ln(B)

Answer by ikleyn(53751)

.
On logarithms, see a bunch of lessons
    - WHAT IS the logarithm
    - Properties of the logarithm
    - Change of Base Formula for logarithms
    - Solving logarithmic equations
    - Using logarithms to solve real world problems
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Logarithms".